OpenStudy (anonymous):

A scientist wants to estimate the volume of Earth and uses a sphere as a model for the planet. According to NASA, Earth’s diameter is 12,756 km. The volume of a sphere is four-thirds of the product of Pi (π) and the radius cubed. Using 3.14 for π and that the diameter of a sphere is equal to two times its radius, estimate the volume of Earth.

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

let me see

OpenStudy (idku):

Diameter is 12,756

OpenStudy (idku):

diameter= radius times 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

25512

OpenStudy (idku):

So you would divide the diameter by 2 to find the radius.

OpenStudy (anonymous):

6378

OpenStudy (idku):

yes, know do you know the formula for the sphere?

OpenStudy (anonymous):

no it does not say

OpenStudy (idku):

\[\frac{4}{3}πr^3---->Sphere\]your π is 3.14, so --->\[\frac{4}{3}\times 3.14 \times r^3\] the radius is 6378\[\frac{4}{3}\times 3.14 \times (6378)^3\] do the 6378^3 first. Plug in and solve.

OpenStudy (anonymous):

ok wow this website really works

OpenStudy (idku):

Sure it does, and hope you understand how to do this on your own in the future. Cuz this is the website's goal- teach.

OpenStudy (anonymous):

ya ok so 6378 x 3

OpenStudy (idku):

6378 is to the third power. then times pi, or times 3.24 then times 4/3 There is your answer.

OpenStudy (anonymous):

i think this is wrong what do u think 20427458

OpenStudy (idku):

1.0862303e+12 I plugged this into google calculator and that's the result.

OpenStudy (anonymous):

sorry im so bad at math ahaha do u want me to do that

OpenStudy (idku):

write this as an answer? I think that when it comes to this, google calc is, although there are things at which it is clearly wrong.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (idku):

Anytime!

OpenStudy (anonymous):

haha would you be available to help with one more

OpenStudy (idku):

Yes, but please post a new question, because threads get too long otherwise; ty!

OpenStudy (anonymous):

ok so i will post a new problem i clicked add a fan

OpenStudy (anonymous):

k im closing it